### v1

# Learning Unified Representation of 3D Gaussian Splatting

Abstract
A well-designed vectorized representation is crucial for the learning systems natively based on 3D Gaussian Splatting. While 3DGS enables efficient and explicit 3D reconstruction, its parameter-based representation remains hard to learn as features, especially for neural-network-based models. Directly feeding raw Gaussian parameters into learning frameworks fails to address the non-unique and heterogeneous nature of the Gaussian parameterization, yielding highly data-dependent models. This challenge motivates us to explore a more principled approach to represent 3D Gaussian Splatting in neural networks that preserves the underlying color and geometric structure while enforcing unique mapping and channel homogeneity. In this paper, we propose an embedding representation of 3DGS based on continuous submanifold fields that encapsulate the intrinsic information of Gaussian primitives, thereby benefiting the learning of 3DGS.
1 Introduction
Recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) have established it as a powerful technique for representing and rendering 3D scenes, enabling high-fidelity, real-time novel view synthesis through explicit parameterization of Gaussian primitives. This representation has catalyzed a growing body of work exploring learning-based methods that operate directly on Gaussian primitives, supporting tasks such as compression (Shin et al., 2025), generation (Yi et al., 2024; Xie et al., 2025), and understanding (Guo et al., 2024). In these pipelines, the native parameterization $\bm{\theta}=\{\bm{\mu},\mathbf{q},\mathbf{s},\mathbf{c},o\}$ is often adopted as the input or output of neural architectures.
Despite its effectiveness in optimization-based reconstruction, we identify fundamental limitations when this parametric representation is employed as a learning space for neural networks. Our analysis reveals three critical issues. First, the mapping from parameters to rendered output exhibits non-uniqueness: multiple distinct parameter configurations yield identical visual results. This occurs through quaternion sign ambiguity (i.e., $\mathbf{q}$ and $\mathbf{q}$ represent identical rotations), symmetry-induced rotation invariance, and equivalences in spherical harmonic coefficients. Second, the parameter components exhibit severe numerical heterogeneity, with scales spanning large magnitude while quaternions remain unit-normalized. Third, these parameters inherently reside on distinct mathematical manifolds: positions in $\mathbb{R}^{3}$, rotations in $\text{SO}(3)$, and appearance in spherical harmonic space, inconsistent with the Euclidean space demands of standard neural architectures.
These theoretical limitations translate into substantial practical failures. Our empirical analysis reveals critical instabilities: negating quaternion values, which should produce equivalent rotations, causes complete decoding failure in parameter-trained autoencoders (see App. Fig. 8). Parametric representations achieve lower reconstruction quality and exhibit severe degradation under domain shift, particularly when transferring from object-level to scene-level data. Moreover, parametric embeddings lack robustness to noise, with minor perturbations causing disproportionate reconstruction errors. These consistent failures indicate fundamental representation deficiencies.
We thus propose a principled alternative: representing each Gaussian primitive as a continuous field defined on its iso-probability surface. This submanifold field representation establishes a unique relationship between Gaussians and their geometric-photometric properties, eliminating the ambiguity inherent in parametric representations. By discretizing this field as a colored point cloud sampled from the probability surface, we obtain a representation that maintains uniform numerical properties and respects the underlying geometric structure. Particularly, we employ a variational autoencoder to learn embeddings from discretized submanifold fields, introducing a Manifold Distance metric based on optimal transport that better aligns with perceptual quality than parameter-space distances. Extensive experiments demonstrate substantial improvements in reconstruction quality, superior cross-domain generalization, and more robust latent representations compared to parametric baselines. The learned embeddings exhibit better noise resilience and semantic structure, validating our submanifold field representation as a more suitable learning space for 3D Gaussian primitives.
To summarize the contribution of this work, we:
  • identify and formally characterize the fundamental limitations of parametric Gaussian representations for neural learning, including non-uniqueness and numerical heterogeneity.
  • propose a submanifold field representation that provides provably unique and geometrically consistent encoding of Gaussian primitives.
  • develop a variational autoencoder framework incorporating a novel Manifold Distance metric based on optimal transport theory for effective learning in the submanifold field representation space with extensive experimental evidence.
Figure 1: A scene of $N$ Gaussian primitives can be represented by NN sets of parameters $\bm{\theta}$ (shown in pink). Data in this parametric space resides on different manifolds and is heterogeneous and non-Euclidean, introducing challenges for encoders to fit disparate data manifolds implicitly. Shown in purple is the proposed representation, instead of relying on Gaussian parameterization, we introduce a canonical submanifold field space $(\mathcal{M},F)$ that uniquely represents a Gaussian primitive with an iso-probability surface.
2 Related Works
3D Gaussian Splatting. Since its re-introduction by Kerbl et al. (2023), 3DGS has rapidly become a core method for novel view synthesis and 3D representation. By placing explicit Gaussian primitives in 3D space and employing efficient rasterization and accumulation, 3DGS achieves real-time rendering with high fidelity (Bao et al., 2025; Lin et al., 2025c). Several studies improve efficiency (Jo et al., 2024; Lee et al., 2024), while others leverage large-scale datasets for generalization (Ma et al., 2025; Li et al., 2025a). At the application level, 3DGS has been adopted for digital human (Li et al., 2024; Kocabas et al., 2024; Wang et al., 2025), self-driving scene modeling (Zhou et al., 2024a; c; Yan et al., 2023), and physics-based simulation (Jiang et al., 2024; Xie et al., 2024; Zhong et al., 2024). Beyond fixed Gaussian parameters, another line augments primitives with latent embeddings to capture semantics, open-vocabulary understanding (Qin et al., 2024) and deformation modeling (Zhobro et al., 2025). These efforts demonstrate that 3DGS provides high-fidelity appearance and serves as a versatile representation with broad potential (Sun et al., 2025).
3DGS Parameters Regression. To enable fast and flexible reconstruction without per-scene optimization, recent works seek to directly obtain Gaussian splats through feedforward prediction networks. For example, Charatan et al. (2024); Chen et al. (2024b) proposed to predict Gaussian parameters directly from multi-view input, while Zheng et al. (2024) generate pixel-wise parameter maps and lift them to 3D via depth estimation. This paradigm has been extended to pose-free settings (Hong et al., 2024; Chen et al., 2024c; Tian et al., 2025), and transformer-based methods further improve scalability and generalization (Li et al., 2025b; Jiang et al., 2025; Lin et al., 2025b).Overall, these methods directly output Gaussian parameters from neural networks, showing that 3DGS can serve as an effective target for network-based prediction in efficient reconstruction.
Embedding Gaussian Primitives. Recent works move beyond reconstruction and encode Gaussian parameters into latent spaces for tasks such as generation, editing, and compression.  Zhou et al. (2024b); Lin et al. (2025a); Wewer et al. (2024) learn structured latent variables from 3D Gaussian space to fulfill generation tasks. Editing and style-transfer methods use diffusion or style conditioning to manipulate Gaussians primitives in latent or rendering spaces (Chen et al., 2024a; Vachha & Haque, 2024; Lee et al., 2025; Palandra et al., 2024; Zhang et al., 2024; Kovács et al., 2024; Yu et al., 2024). Other works improve rendering quality by optimizing Gaussian parameters under diffusion priors (Tang et al., 2023; Yi et al., 2024; Chen et al., 2024d), while compression methods (Girish et al., 2024; Yang et al., 2025) reduce storage and computation by quantizing and embedding Gaussian parameters. These approaches show the potential of embedding Gaussians into neural latent spaces, but they assume Gaussian parameters are naturally compatible with neural learning, overlooking that these parameters were designed for optimization-based reconstruction. This oversight underlies our analysis in Section 3 and our proposal of a more suitable formulation.
3 Method
3.1 Preliminaries: Gaussian Splatting Parameterization
A scene under 3D Gaussian Splatting is represented as a set of $N$ oriented, and view-dependently colored Gaussian primitives $\{\mathcal{G}_{i}\}_{i=1}^{N}$, each contributing to the final rendered image via rasterization and alpha compositing. Each Gaussian primitive $\mathcal{G}_{i}$ is usually represented by a parameter tuple $\bm{\theta}_{i}=\{\bm{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\mathbf{c}_{i},o_{i}\}$, where:
  • $\bm{\mu}_{i}\in\mathbb{R}^{3}$: the center position of the Gaussian in world coordinates;
  • $\mathbf{q}_{i}\in\text{SO}(3)$: a unit quaternion representing the local rotation;
  • $\mathbf{s}_{i}\in(\mathbb{R}^{+})^{3}$: scale parameters along the rotated axes;
  • $\mathbf{c}_{i}\in\mathbb{R}^{3\times K}$: spherical harmonic (SH) coefficients for view-dependent color for $K\in\mathbb{Z}$;
  • $o_{i}\in\mathbb{R}$: a logit-transformed opacity value $\alpha_{i}=\sigma(o_{i})$, where $\sigma$ is a sigmoid function.
The local geometry of the Gaussian is governed by its covariance matrix, constructed as
$\Sigma_{i}=R(\mathbf{q}_{i})\,\text{diag}(\mathbf{s}_{i})^{2}\,R(\mathbf{q}_{i})^{\top},$ (1)
where $R(\mathbf{q}_{i})$ is the rotation matrix corresponding to the quaternion $\mathbf{q}_{i}$. This defines an ellipsoidal spatial density, centered at $\bm{\mu}_{i}$, whose shape and orientation determine the contribution of $\mathcal{G}_{i}$ to the rendered scene. The color at a given view direction $\mathbf{d}\in\mathbb{S}^{2}$ is computed per channel using SH basis functions denoted by
$\text{Color}_{i}(\mathbf{d})=\left[\text{SH}^{r}_{i}(\mathbf{d}),\text{SH}^{g}_{i}(\mathbf{d}),\text{SH}^{b}_{i}(\mathbf{d})\right]^{\top},$ (2)
where $\text{SH}^{c}_{i}(\mathbf{d})$ in $c$-channel is calculated by $\sum_{l=0}^{L_{\text{max}}}\sum_{m=-l}^{l}(\mathbf{c}_{i})_{c,(l,m)}\cdot Y_{l}^{m}(\mathbf{d})$ $Y_{l}^{m}$ is the real-valued spherical harmonic of degree $l$ and order $m$. The final rendering aggregates contributions from all $\mathcal{G}_{i}$ via a soft visibility-weighted compositing process. This native parameterization is well-suited for gradient-based scene optimization. However, it introduces significant challenges when used as a representation for learning.
3.2 Parameterization is Ill-Suited as a Learning Space
The parameter representation $\bm{\theta}$ poses fundamental challenges when used as a learning space for neural networks. We identify two critical issues: representation non-uniqueness and numerical heterogeneity. Each undermines the stability and effectiveness of neural network training.
Representation Non-uniqueness. The parametric representation suffers from a many-to-one mapping that violates basic requirements for stable learning. To understand this, we first formalize what rendering effect a single Gaussian primitive produces.
Definition 1 (Single Gaussian Radiance Field (SGRF))
A SGRF is a radiance field $\phi:\mathbb{R}^{3}\times\mathbb{S}^{2}\rightarrow{\mathbb{R}}^{3}$, the field is defined by the local density at point $\mathbf{x}\in\mathbb{R}^{3}$ along direction $\mathbf{d}\in\mathbb{S}^{2}$:
$\phi_{\mathcal{G}}(\mathbf{x},\mathbf{d})=\rho_{\mathcal{G}}(\mathbf{x})\cdot c_{\mathcal{G}}(\mathbf{d}),$ (3)
where $\rho_{\mathcal{G}}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\Sigma^{-1}(\mathbf{x}-\bm{\mu})\right)$ is a volume density function and $c_{\mathcal{G}}(\mathbf{d})$ is a color radiance field coupled with opacity. Specifically, given a parameter set $\bm{\theta}=\{\bm{\mu},\mathbf{q},\mathbf{s},\mathbf{c},o\}$, $\Sigma$ can be derived by  equation 1 and $c_{\mathcal{G}}(\mathbf{d})=\sigma(o)\cdot\operatorname{Color}(\mathbf{d})$ can be derived by  equation 2.
The SGRF, derived from the multi-Gaussian rendering framework by Kerbl et al. (2023), specifies how the final value at any pixel is rendered in a scene containing only one Gaussian splat. Furthermore, let $\Phi$ be the space of SGRFs, and $\Theta\subseteq\mathbb{R}^{|\bm{\theta}|}$ be the paramater space of Gaussian primitives, each parameter set $\bm{\theta}\in\Theta$ provides a complete representation that generates a correponding field $\phi_{\mathcal{G}}\in\Phi$. We indicate that a single SGRF may correspond to multiple parameterizations of Gaussian primitives, as formalized in the following proposition.
Proposition 1 (Non-uniqueness of the SGRF Parametric Representation)
The parametric representation of a SGRF is not unique. Formally, there exist at least two distinct parameter sets, $\bm{\theta}_{1}\in\Theta$ and $\bm{\theta}_{2}\in\Theta$ with $\bm{\theta}_{1}\neq\bm{\theta}_{2}$, that generate the exact same field $\phi_{\mathcal{G}}\in\Phi$.
The non-uniqueness is from quaternion sign ambiguity, geometric symmetries, and rotation-spherical harmonic interactions producing equivalent parameter combinations (see Appendix for proof). The non-uniqueness of $\bm{\theta}$ will create “embedding collisions” where different parameter vectors produce identical rendered output (Wang & Isola, 2020). This makes the learning objective $\|\bm{\theta}_{\text{pred}}-\bm{\theta}_{\text{target}}\|_{p}$ ambiguous, as multiple parameter configurations can achieve the same visual result. The resulting conflicting gradients lead to training instability and poor convergence indicated by Bengio et al. (2013).
Numerical Heterogeneity The parameter components violate the homogeneous distribution assumption of standard neural architectures. Neural networks typically assume features share similar statistical properties for effective gradient flow (Ioffe & Szegedy, 2015). However, 3D Gaussian parameters span vastly different ranges. For example, pre-activation scales can range from 15-15 to $3$, while quaternions stay unit-normalized. More fundamentally, these parameters follow different distributions and live on different manifolds: positions $\bm{\mu}\in\mathbb{R}^{3}$, rotations $\mathbf{q}\in\text{SO}(3)$, scales $\mathbf{s}\in(\mathbb{R}^{+})^{3}$, and SH coefficients $c$ with exponential decay. Concatenating them ignores their heterogeneous nature. Small noises in quaternions can drastically alter geometry, while small noise in high-order SH coefficients is negligible, yet the network treats all dimensions equally.
The non-uniqueness and numerical heterogeneity of the native parameter space $\bm{\theta}$ make it unsuitable for neural network learning, which would generate unstable embeddings (see our experiments in Sec. 4.4 and App. D). We therefore introduce a submanifold field representation that ensures unique mappings and respects the geometric structure of 3D Gaussians.
3.3 Representation on Submanifold Field
To address this issue, we propose converting each Gaussian primitive $\mathcal{G}_{i}$ to a novel geometric representation $\mathcal{E}_{i}$, which is a color field defined on a 2D submanifold in 3D Euclidean space, as illustrated in Fig. 1. For a Gaussian density $\mathcal{N}(\mathbf{x};\bm{\mu}_{i},\Sigma_{i})$, we define the iso-probability surface at fixed radius $r$ as:
$\mathcal{M}_{i}=\left\{\mathbf{x}\in\mathbb{R}^{3}\mid(\mathbf{x}-\bm{\mu}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{x}-\bm{\mu}_{i})=r^{2}\right\},$ (4)
which forms an ellipsoid surface, namely, a two-dimensional submanifold, centered at $\bm{\mu}_{i}$. On this submanifold, we define a field function:
$F_{i}(\mathbf{x})=\sigma(o_{i})\cdot\text{Color}_{i}(\mathbf{d}_{\mathbf{x}}),$ (5)
where $\mathbf{d}_{\mathbf{x}}=(\mathbf{x}-\bm{\mu}_{i})/\|\mathbf{x}-\bm{\mu}_{i}\|$ denotes the unit direction vector for $\mathbf{x}\in\mathcal{M}_{i}$, and $\text{Color}_{i}(\cdot)$ represents the view-dependent color parameterization as in equation 2. Let $\mathbb{M}$ be the space of all possible iso-probability submanifolds as defined in  equation 4, we define our unified representation space as:
$\mathscr{E}=\left\{\mathcal{E}_{i}=(\mathcal{M}_{i},F_{i})\mid\mathcal{M}_{i}\in\mathbb{M},\,F_{i}:\mathcal{M}_{i}\rightarrow\mathbb{R}^{3}\right\},$ (6)
The representation $\mathcal{E}_{i}\in\mathscr{E}$ encodes both geometric properties (shape, orientation) via $\mathcal{M}_{i}$ and appearance attributes (view-dependent color) via $F_{i}$ in a continuous framework. We have the following proposition (proof is provided in App. B).
Proposition 2 (Uniqueness of Submanifold Field Representation)
For every SGRF $\phi_{\mathcal{G}}\in\Phi$, there exists a unique corresponding representation $\mathcal{E}\in\mathscr{E}$. This establishes a one-to-one correspondence between the elements of $\Phi$ and $\mathscr{E}$ . Formally, for any two distinct fields $\phi_{\mathcal{G},1},\phi_{\mathcal{G},2}\in\Phi$, their corresponding representations $\mathcal{E}_{1},\mathcal{E}_{2}\in\mathscr{E}$ are also distinct.
The submanifold field $\mathcal{E}_{i}$ thus provides a numerically stable and provably unique representation space on which we can safely build learning objectives and neural architectures.
3.4 Encode Submanifold Fields as Embeddings
Figure 2: To embed the proposed submanifold field representation into a vector form suitable for neural networks, we devise a submanifold field variational auto-encoder (SF-VAE) that embeds any input submanifold field as a 32-D vector, then reconstructs the original parameter set $\bm{\theta}_{i}$. SF-VAE learns in our new representation space instead of the parametric space.
We design a variational auto-encoder to encode submanifold field representation, shown in Fig. 2. The network architecture, learning objectives and dataset are introduced.
Encoder-decoder Architecture. We employ a point-cloud-based network to encode and decode one sub-manifold field. Particularly, we uniformly sample $P$ points from the submanifold field $\mathcal{E}$ as a colored point cloud $\mathcal{P}=\bigl\{(\mathbf{x}_{m},F(\mathbf{d}_{\mathbf{x}_{m}}))\bigr\}_{m=1}^{P}$. We then employ a PointNet (Qi et al., 2017) encoder $f$ to obtain latent embedding by $\mathbf{z}\sim f(\mathbf{z}\mid\mathcal{P})$ where $\mathbf{z}\in\mathbb{R}^{D}$ is the embedding with dimension $D$. The decoder $g$ consists of two neural networks, namely, the coordinates transform network $g_{c}:\mathbb{R}^{3}\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{3}$ and color field $g_{f}:\mathbb{R}^{3}\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{3}$. The decoded point cloud from decoder is given by,
$\hat{\mathcal{P}}=g(\mathbf{z},\mathcal{U}_{P^{\prime}})=\{g_{c}([\mathbf{e}_{n},\mathbf{z}]),g_{f}(\left[g_{c}([\mathbf{e}_{n},\mathbf{z}]),\mathbf{z}\right])\}_{n=1}^{P^{\prime}}$ (7)
where $\mathcal{U}_{P^{\prime}}=\{\mathbf{e}_{n}\}_{n=1}^{P^{\prime}}$ is a set of coordinates sampled from a unit sphere surface. Such canonical set works as the initial input for two implicit functions $g_{c}$ and $g_{f}$, and queries new coordinates and color field. Furthermore, to recover the original Gaussian parameters $\bm{\theta}_{i}$ for rendering purposes, we estimate the covariance matrix $\Sigma_{i}$ by principal component analysis (PCA), and SH coefficients $\mathbf{c}_{i}$ by fitting the spherical harmonics to $\hat{\mathcal{P}}$.
[…]
$W_{2}^{2}(\mathcal{E},\hat{\mathcal{E}})=\inf_{\gamma\in\Gamma(\hat{\sigma},\hat{\sigma}^{\prime})}\int_{\mathcal{M}\times\hat{\mathcal{M}}}d^{2}\big((\mathbf{x},c_{x}),(\mathbf{y},c_{y})\big)d\gamma(\mathbf{x},\mathbf{y})$ (8)
where $c_{x}=F(\mathbf{d}_{\mathbf{x}})$, $c_{y}=\hat{F}(\mathbf{d}_{\mathbf{y}})$, $\Gamma(\hat{\sigma}_{i},\hat{\sigma}_{j})$ is the set of all joint probability measures (transport plans) with marginals $\hat{\sigma}_{i}$ and $\hat{\sigma}_{j}$, and the ground distance is defined as
$d^{2}\big((\mathbf{x},c_{x}),(\mathbf{y},c_{y})\big)=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}+\lambda\|c_{x}-c_{y}\|_{2}^{2},$ (9)
with $\lambda\in\mathbb{R}^{+}$ balancing spatial and color terms. In practice, both $\mathcal{M}$ and $\hat{\mathcal{M}}$ are discretized as colored point clouds $\mathcal{P}$ and $\hat{\mathcal{P}}$. The empirical Wasserstein-2 distance $\hat{W}$ is then computed between these point clouds by
$\hat{W}_{2}^{2}(\mathcal{P},\hat{\mathcal{P}})=\min_{\mathbf{\Gamma}\in\Gamma(\hat{\sigma},\hat{\sigma}^{\prime})}\sum_{(\mathbf{x}_{i},c_{x_{i}})\in\mathcal{P}}\sum_{(\mathbf{y}_{j},c_{y_{j}})\in\hat{\mathcal{P}}}\mathbf{\Gamma}_{ij}\left(d^{2}\big((\mathbf{x}_{i},c_{x_{i}}),(\mathbf{y}_{j},c_{y_{j}})\big)\right).$ (10)
Finally, the learning objective for variational auto-encoder is
$\mathcal{L}_{\text{VAE}}=\mathbb{E}_{\hat{\mathcal{P}}\sim\operatorname{VAE}(\mathcal{P})}\left(\hat{W}_{2}^{2}(\mathcal{P},\hat{\mathcal{P}})+\beta\cdot d_{\text{KL}}\left(f(\mathbf{z}\mid\mathcal{P})\|\mathcal{N}(0,\mathbf{I})\right)\right),$ (11)
where $\operatorname{VAE}(\mathcal{P})=g(f(\mathbf{z}\mid\mathcal{P}),\mathcal{U}_{P^{\prime}})$ and the second term is the KL divergence loss for variational auto-encoder implementation, and $\beta$ is a balance factor.
Dataset Preparation. Since this embedding model only encodes single Gaussian primitives, which have no semantic meaning out of a scene’s global context, we can use a randomly generated dataset of Gaussian primitives to train this model, thus making it domain-invariant to data. The implementation details of this generated dataset can be found in App. C.
4 Experiments
4.1 Evaluation Setup
Baseline Implementation Details. To isolate the effect of representation choice, we adopt a self-implemented encoder–decoder framework for both the parametric representation $\bm{\theta}$ and the proposed submanifold field representation $\mathcal{E}$. While comparisons with existing 3DGS learning methods are possible, they typically involve task-specific architectures that confound the role of representation itself. Direct reuse of prior pipelines would not yield a controlled comparison, so we implement both baselines in the same VAE-style framework to attribute differences solely to the representation.
We implement and train three size-matched embedding models: our submanifold field VAE (Sec. 3.4), and two baseline parametric VAEs operating directly on $\bm{\theta}$. For the parametric models, each Gaussian primitive is represented as a 56-D vector (3+4+3K+13{+}4{+}3K{+}1 for $L_{\mathrm{max}}{=}3$), omitting global coordinates to match the SF-VAE setting. A three-layer MLP encodes this input to a 32-D latent ($56\!\!\rightarrow\!\!512\!\!\rightarrow\!\!512\!\!\rightarrow\!\!32\times 2$), and the decoder, either uses a MLP to map the latent back to $\widehat{\bm{\theta}}_{i}$, or uses the same decoder of SF-VAE to map to $\hat{\mathcal{P}}$. This setting further decouples evaluation results with the training objective functions. Apart from input dimension, all models share identical depth, width, latent size (32), and optimizer settings, ensuring a matched capacity.
Datasets. We evaluate the proposed representation and compare it with the baseline primarily using two datasets. For object-level tasks, we utilize ShapeSplat (Ma et al., 2025), a large-scale 3DGS dataset derived from ShapeNet (Chang et al., 2015), comprising 52K objects across 55 categories. For scene-level experiments, we employ Mip-NeRF 360 (Barron et al., 2022), which contains 7 medium-scale scenes with abundant high-frequency details. Additionally, unless stated otherwise, we train the embedding models using the randomly generated Gaussian primitive dataset, with 500K randomly generated data samples; implementation details are provided in App. C.
Evaluation Metrics. To comprehensively assess both perceptual fidelity and representation quality, we report PSNR, SSIM, and LPIPS on rasterized reconstructions against ground truth Gaussian splats, as well as $L_{1}$ distance in the Gaussian parameter space. Crucially, we also include our proposed Manifold Distance (M-Dist) as an evaluation criterion. By cross-comparing M-Dist with parameter-space distances ($L_{1}$$L_{2}$), we can demonstrate that M-Dist aligns more closely with perceptual “gold standard” metrics such as PSNR and LPIPS, validating our claims in Sec. 3
4.2 Quantitative Reconstruction Results
Table 1: Reconstruction quality comparison for object-level (ShapeSplat) and scene-level (Mip-NeRF 360) datasets. All models trained on the randomly generated dataset. The three models have a parameter count of 0.62M, 0.66M and 0.62M respectively. The relatively extreme perceptual metrics values in ShapeSplat come from the use of background during measurement.
Input Representation Encoder Decoder PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ M-Dist $L_{1}$-Dist
[…]
Parametric MLP MLP 37.512 0.888 0.152 0.184 0.040
Parametric MLP SF-VAE 44.725 0.896 0.136 0.051 0.097
Submanifold Field SF-VAE SF-VAE 63.408 0.990 0.010 0.041 0.098
[…]
Parametric MLP MLP 18.818 0.564 0.452 0.510 0.034
Parametric MLP SF-VAE 20.923 0.730 0.359 0.055 0.173
Submanifold Field SF-VAE SF-VAE 29.833 0.953 0.079 0.048 0.179
We present a comprehensive quantitative and qualitative analysis of reconstruction quality for both object-level and scene-level data, as summarized in Tab. 4.2 and Fig. 3. All models are trained on the same randomly generated 3D Gaussian primitives dataset and evaluated on ShapeSplat and Mip-NeRF 360, using three matched encoder-decoder configurations to control for bias. Across all perceptual metrics (PSNR, SSIM, LPIPS), the submanifold field representation consistently outperforms parametric baselines. For example, on ShapeSplat, SF-VAE achieves substantially higher PSNR and SSIM and a much lower LPIPS, indicating both improved fidelity and perceptual quality. Similar performance gains are observed in scene-level reconstruction, where the submanifold field model demonstrates better performance across diverse spatial contexts.
Importantly, the Manifold Distance (M-Dist) metric shows a stronger empirical correlation with perceptual metrics like PSNR and LPIPS than traditional $L_{1}$ parameter distances, supporting our claim that M-Dist is a more robust and meaningful similarity measure for 3D Gaussian representations, truthfully reflecting perceptual differences rather than merely parameter discrepancies. The consistent improvement margin across both datasets highlights the advantage of learning in the submanifold field space, which better preserves intrinsic structure and view-dependent appearance, confirming the efficacy of our representation for high-fidelity 3D Gaussian modeling.
Figure 3: Qualitative results for rasterized reconstruction. Samples selected arbitrarily from Mip-NeRF 360 and ShapeSplat. Parametric models can induce confusion in rotation space, failing to predict the correct rotation.
4.3 Cross-Domain Generalization
Table 2: Results on representation generalization: train on one dataset, test on the other.
Train set Test set Input Represent. Encoder / Decoder PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ShapeSplat Mip-NeRF. Parametric MLP / MLP 9.753 0.356 0.615
Parametric MLP / SF-VAE 14.845 0.675 0.336
Submanifold Field SF-VAE / SF-VAE 19.194 0.821 0.309
Mip-NeRF. ShapeSplat Parametric MLP / MLP 55.624 0.957 0.067
Parametric MLP / SF-VAE 60.777 0.987 0.013
Submanifold Field SF-VAE / SF-VAE 62.576 0.990 0.014
We evaluate cross-domain generalization by training on one dataset and testing on the other (object-level $\leftrightarrow$ scene-level) under an identical training protocol and capacity budget as in the reconstruction study. Concretely, we train either the proposed SF-VAE or the parametric MLP baselines on a source set and evaluate zero-shot on a target set, rendering novel views and reporting PSNR, SSIM, and LPIPS averaged over test samples (see Tab.  2).
Across both transfer directions, the submanifold field model consistently achieves higher reconstruction quality than the parametric baseline, with the largest margins when transferring from object-level sources to larger-scale scenes, indicating reduced sensitivity to dataset-specific statistics (scale, lighting, SH complexity).
4.4 Latent Space Evaluation
Robustness to Noise. To evaluate the submanifold field embedding’s robustness to noise, we gradually add higher levels of gaussian noise to the embedding space of each representation and test their reconstruction quality and M-Dist. As shown in Fig. 4, the embedding space of submanifold field model is more robust to random perturbation, this makes submanifold field embeddings a better learning target since it is less sensitive to potential systematic noise.
Figure 4: Parametric results use MLP encoder and SF-VAE’s decoder. Left: Sample visual comparison on robustness to noise in embedding space. Right: Comparison on M-Dist for different noise levels added to embedding space, tested on Mip-NeRF 360.
Latent Space Clustering. To further probe the semantic structure of the learned embedding spaces, we perform unsupervised graph clustering on both the raw Gaussian parameter space and the embedding outputs of each model. As visualized in Fig. 5, clusters formed in the submanifold field embedding space exhibit more detailed semantic separation against the reference images compared to those formed using normalized parameters or parametric embeddings. For example, SF-VAE’s embedding clustering in the first line of Fig. 5 outlines clearer separation of foreground objects with the background. The clusters appear smoother, less noisy, and with clearer boundaries, showing an ability to distinguish between different entities. This indicates that the submanifold field embedding captures more dense semantics and discriminative features, validating its usefulness.
Figure 5: Qualitative evaluation of unsupervised graph clustering, compared using Gaussian parameters and the embedding of the respective models, using the same hyperparameters. Submanifold field embeddings show better preservation of detailed semantics, validating its usefulness.
Latent Space Interpolation. To evaluate the regularity of the latent space of the proposed representation, we randomly sample pairs of source and target Gaussian primitives $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ and linearly interpolate each pair for a fixed number of steps $n=7$. Compared with parametric space, the interpolation in submanifold field embedding space shows a smooth transition path, while interpolation in parametric space shows undesired jitter in rotation and scale, indicating space irregularities, see App. D. This highlights the motivation to learn in the unified submanifold field embedding space.
4.5 More Studies
Figure 6: Ablation study results, tested on Mip-NeRF 360. From left to right: (a) embedding space dimension, (b) generated training dataset size, (c) Submanifold Field sample size (all values normalized to $[0,1]$, with LPIPS inversely normalized).
Latent Space Dimension. We evaluated different embedding space dimensions for the SF-VAE model to meet the best trade-off between compression and reconstruction quality. All models are trained on the generated dataset with $L=3$ order Spherical Harmonics. Results shown in Fig. 6 (a), 32 is the optimal balance point between reconstruction quality and latent space compression. All values are tested with baseline input/output of $P=12^{2}$. While this work does not specifically focus on compression effectiveness, embedding space robustness shown in Sec. 4.4 suggests potential in further latent tokenization and quantization.
Size of Training Set. To empirically determine the number of random training samples required to achieve the best reconstruction results, we tested a wide spectrum of training sample sizes, ranging from 5K to 500K (baseline). The results indicate the proposed representation is data efficient. When only using 2% of the baseline training sample, our model can achieve close-to-baseline performance.
Point Sample Size of Submanifold Fields. To ensure the submanifold fields are truthfully represented in a discrete manner, we evaluate different sample sizes $P$. Going from the lowest tested $P=6^{2}$ to the baseline $P=12^{2}$, we observe a gradual improvement in reconstruction quality, while going above $P=12^{2}$ yields very little improvement. Since $P$ directly correlates to the computational efficiency of the submanifold field model (see below), we keep $P=12^{2}$.
Computational Efficiency. Increasing the point sample size $P$ leads to higher compute time and memory consumption. Encoding time stays low since a lightweight PointNet-based encoder shares weights with all input points, ensuring efficient encoding of large-scale scenes, with an inference speed of 1.72s per 1 million Gaussians for $P=12^{2}$ using a batch size of 4096 on an RTX 5090. Query time during decoding is $O(P)$ which is $O(n^{2})$ w.r.t. the grid size $n$. When the computation is not bottlenecked and uses a large enough batch size, the amortized time complexity is $\Theta(n)$.
5 Conclusion and Limitations
We introduced a geometry-aware submanifold field representation for 3D Gaussian Splatting that maps each primitive to a color field on a canonical iso–probability ellipsoid and proved the mapping is injective over core attributes. Built on this representation, our SF–VAE learns semantically meaningful latents and yields higher-fidelity reconstructions and stronger zero-shot generalization than capacity-matched raw-parameter baselines; our manifold distance (M-Dist) further aligns training and evaluation with geometric/perceptual similarity.
Limitations and Outlook. Our current setup operates at the single-Gaussian level, while this ensures data invariance, it omits explicit inter-splat structure modeling for more complex representation learning. Promising directions include set/scene-level encoders with permutation-invariant attention, point cloud to 3DGS inpainting, generative modeling with submanifold field embeddings, temporal extensions to dynamic scenes, and applications to compression, retrieval, and regularization in broader 3DGS pipelines.
Appendix A Proof of Proposition 1
We construct two distinct parameter sets, $\bm{\theta}_{1}=\{\mathbf{q}_{1},\mathbf{s}_{1},\mathbf{c}_{1},o_{1}\}$ and $\bm{\theta}_{2}=\{\mathbf{q}_{2},\mathbf{s}_{2},\mathbf{c}_{2},o_{2}\}$ with $\bm{\theta}_{1}\neq\bm{\theta}_{2}$, that generate the identical Single Gaussian Radiance Field (SGRF). The identity $L_{\mathcal{G}_{1}}(\mathbf{x},\mathbf{d})=L_{\mathcal{G}_{2}}(\mathbf{x},\mathbf{d})$ for all $(\mathbf{x},\mathbf{d})$ requires two conditions to be met:
  1. 1.
    Geometric Equivalence: The covariance matrices must be equal, $\bm{\Sigma}_{1}=\bm{\Sigma}_{2}$, which implies that the volume densities $\sigma_{\mathcal{G}}(\mathbf{x})$ are identical. We also assume equal opacity, $o_{1}=o_{2}$.
  2. 2.
    Appearance Equivalence: The view-dependent color functions must be identical, which requires $\text{SH}(\mathbf{c}_{1},\mathbf{R}_{1}^{\top}\mathbf{d})=\text{SH}(\mathbf{c}_{2},\mathbf{R}_{2}^{\top}\mathbf{d})$ for all $\mathbf{d}\in\mathbb{S}^{2}$, where $\mathbf{R}$ is the rotation matrix for the quaternion $\mathbf{q}$.
We construct a distinct parameter set $\bm{\theta}_{2}$ by considering a discrete symmetry of the Gaussian ellipsoid. Let $\bm{\theta}_{1}=\{\mathbf{q}_{1},\mathbf{s}_{1},\mathbf{c}_{1},o_{1}\}$ be an initial parameterization.
Let $\mathbf{R}_{\text{flip}}$ be a rotation matrix corresponding to a 180-degree rotation about one of the local axes (e.g., the z-axis), such that $\mathbf{R}_{\text{flip}}=\text{diag}(-1,-1,1)$. $\mathbf{R}_{\text{flip}}$ is its own inverse, $\mathbf{R}_{\text{flip}}^{\top}=\mathbf{R}_{\text{flip}}$. We define a new parameter set $\bm{\theta}_{2}$ as follows:
  • Let the new rotation be $\mathbf{R}_{2}=\mathbf{R}_{1}\mathbf{R}_{\text{flip}}$. This defines a new quaternion $\mathbf{q}_{2}\neq\mathbf{q}_{1}$.
  • Let the scales and opacity remain unchanged: $\mathbf{s}_{2}=\mathbf{s}_{1}$ and $o_{2}=o_{1}$.
First, we verify geometric equivalence. The new covariance matrix $\bm{\Sigma}_{2}$ is:
$\bm{\Sigma}_{2}$ $=\mathbf{R}_{2}\text{diag}(\mathbf{s}_{2})^{2}\mathbf{R}_{2}^{\top}=(\mathbf{R}_{1}\mathbf{R}_{\text{flip}})\text{diag}(\mathbf{s}_{1})^{2}(\mathbf{R}_{1}\mathbf{R}_{\text{flip}})^{\top}$
$=\mathbf{R}_{1}\left(\mathbf{R}_{\text{flip}}\text{diag}(\mathbf{s}_{1})^{2}\mathbf{R}_{\text{flip}}^{\top}\right)\mathbf{R}_{1}^{\top}$
Since $\mathbf{R}_{\text{flip}}$ is diagonal, it commutes with the diagonal scaling matrix $\text{diag}(\mathbf{s}_{1})^{2}$, meaning the term in parentheses equals $\text{diag}(\mathbf{s}_{1})^{2}$. Therefore,
$\bm{\Sigma}_{2}=\mathbf{R}_{1}\text{diag}(\mathbf{s}_{1})^{2}\mathbf{R}_{1}^{\top}=\bm{\Sigma}_{1}.$
Geometric equivalence is satisfied.
Next, for appearance equivalence, we must find SH coefficients $\mathbf{c}_{2}$ such that:
$\text{SH}(\mathbf{c}_{1},\mathbf{R}_{1}^{\top}\mathbf{d})=\text{SH}(\mathbf{c}_{2},(\mathbf{R}_{1}\mathbf{R}_{\text{flip}})^{\top}\mathbf{d})=\text{SH}(\mathbf{c}_{2},\mathbf{R}_{\text{flip}}^{\top}\mathbf{R}_{1}^{\top}\mathbf{d}).$
Let $\mathbf{v}=\mathbf{R}_{1}^{\top}\mathbf{d}$ be the view direction in the local frame of the first Gaussian. The condition becomes $\text{SH}(\mathbf{c}_{1},\mathbf{v})=\text{SH}(\mathbf{c}_{2},\mathbf{R}_{\text{flip}}^{\top}\mathbf{v})$. This states that the function defined by $\mathbf{c}_{2}$ when evaluated on a transformed vector must equal the function defined by $\mathbf{c}_{1}$ on the original vector. This is equivalent to stating that the function itself has been rotated. The properties of spherical harmonics guarantee that for any rotation, there exists a linear transformation (the Wigner D-matrix $\mathbf{D}$) that maps the original coefficients to the new ones. We can therefore find $\mathbf{c}_{2}$ such that:
$\mathbf{c}_{2}=\mathbf{D}(\mathbf{R}_{\text{flip}})\mathbf{c}_{1}.$
We have constructed a parameter set $\bm{\theta}_{2}=\{\mathbf{q}_{2},\mathbf{s}_{1},\mathbf{c}_{2},o_{1}\}$, which is distinct from $\bm{\theta}_{1}$ (since $\mathbf{q}_{2}\neq\mathbf{q}_{1}$) yet defines the identical radiance field. This discrete symmetry exists for any Gaussian, proving the general proposition.
Furthermore, this non-uniqueness expands from a discrete set to a continuous manifold of solutions for symmetric geometries.
  • Case A (Isotropic): If $\mathbf{s}=(s,s,s)^{\top}$, the covariance matrix $\bm{\Sigma}=s^{2}\mathbf{I}$ is invariant under any rotation. This gives rise to a continuous, multi-parameter family of equivalent solutions.
  • Case B (Spheroidal): If two scale components are equal (e.g., $\mathbf{s}=(s_{a},s_{a},s_{b})^{\top}$), $\bm{\Sigma}$ is invariant to any rotation around the local axis of symmetry, resulting in a one-parameter continuous family of redundant solutions.
In all such cases, a corresponding transformation on the SH coefficients preserves appearance equivalence. Since non-unique parameterizations exist in all cases, the proposition is proven.
Appendix B Proof of Uniqueness of the Representation $\mathcal{E}_{i}=(\mathcal{M}_{i},F_{i})$
Assume $\phi_{\mathcal{G}_{1}}=\phi_{\mathcal{G}_{2}}$, namely, $\rho_{\mathcal{G}_{1}}(\mathbf{x})=\rho_{\mathcal{G}_{2}}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{3}$ and $c_{\mathcal{G}_{1}}(\mathbf{d})=c_{\mathcal{G}_{2}}(\mathbf{d}),\forall\mathbf{d}\in\mathbb{S}^{2}$. We show this leads to $\mathcal{E}_{1}=\mathcal{E}_{2}$ for two SGRFs. This implies both $\mathcal{M}_{1}=\mathcal{M}_{2}$ and $F_{1}=F_{2}$.
The volume densities are equal:
$\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu}_{1})^{\top}\mathbf{\Sigma}_{1}^{-1}(\mathbf{x}-\bm{\mu}_{1})\right)=\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu}_{2})^{\top}\mathbf{\Sigma}_{2}^{-1}(\mathbf{x}-\bm{\mu}_{2})\right)$
Taking the natural logarithm of both sides yields that the quadratic forms are identical for all $\mathbf{x}$:
$(\mathbf{x}-\bm{\mu}_{1})^{\top}\mathbf{\Sigma}_{1}^{-1}(\mathbf{x}-\bm{\mu}_{1})=(\mathbf{x}-\bm{\mu}_{2})^{\top}\mathbf{\Sigma}_{2}^{-1}(\mathbf{x}-\bm{\mu}_{2})$
An unnormalized Gaussian distribution is uniquely defined by its mean and covariance. Therefore, this equality implies $\bm{\mu}_{1}=\bm{\mu}_{2}$ and $\mathbf{\Sigma}_{1}=\mathbf{\Sigma}_{2}$.
The submanifold $\mathcal{M}_{i}$ is defined as the level set:
$\mathcal{M}_{i}=\{\mathbf{x}\in\mathbb{R}^{3}\mid(\mathbf{x}-\bm{\mu}_{i})^{\top}\mathbf{\Sigma}_{i}^{-1}(\mathbf{x}-\bm{\mu}_{i})=r^{2}\}$
Since the parameters $(\bm{\mu}_{i},\mathbf{\Sigma}_{i})$ that define the level set are identical for both primitives, the resulting sets of points must also be identical. Thus, $\mathcal{M}_{1}=\mathcal{M}_{2}$.
The submanifold color field $F_{i}$ is defined for a point $\mathbf{x}\in\mathcal{M}_{i}$ as:
$F_{i}(\mathbf{x})=c_{\mathcal{G}_{i}}(\mathbf{d}_{\mathbf{x}}),\quad\text{where}\quad\mathbf{d}_{\mathbf{x}}=(\mathbf{x}-\bm{\mu}_{i})/\lVert\mathbf{x}-\bm{\mu}_{i}\rVert$
From the hypothesis, we know that $c_{\mathcal{G}_{1}}(\mathbf{d})=c_{\mathcal{G}_{2}}(\mathbf{d})$ holds for any unit direction vector $\mathbf{d}\in\mathbb{S}^{2}$.
For any point $\mathbf{x}$ on the common manifold $\mathcal{M}=\mathcal{M}_{1}=\mathcal{M}_{2}$, its corresponding direction vector $\mathbf{d}_{\mathbf{x}}$ is an element of $\mathbb{S}^{2}$. We can therefore apply the hypothesis for this specific direction:
$c_{\mathcal{G}_{1}}(\mathbf{d}_{\mathbf{x}})=c_{\mathcal{G}_{2}}(\mathbf{d}_{\mathbf{x}})$
By the definition of $F_{i}$, this directly implies:
$F_{1}(\mathbf{x})=F_{2}(\mathbf{x})$
This holds for all $\mathbf{x}\in\mathcal{M}$. Thus, the color fields $F_{1}$ and $F_{2}$ are identical.
Appendix C Implementation details
C.1 Submanifold Field VAE
For completeness, we note a few aspects not detailed in the main text. Also see Alg. 1 for steps in one training step.
Uniform Point Sampling. The submanifold field $\mathcal{E}_{i}$ is discretized by using a uniform mesh grid of size $(n,n)$ to sample $P=n^{2}$ points on the ellipsoidal surface $\mathcal{M}_{i}$ with respect to area, forming $\mathcal{P}_{i}=\{(\mathbf{x}_{i,k},F_{i}(\mathbf{x}_{i,k}),\alpha_{i})\}_{k=1}^{P}$.
Decoding Gaussian Parameters. After decoding, we recover Gaussian parameters from the reconstructed point cloud by first applying batched PCA to estimate the ellipsoid axes and orientation: we compute the mean and covariance of the points, perform eigen decomposition to obtain principal axes, and ensure a right-handed coordinate system. The logarithm of the axis lengths gives the scale parameters, and the rotation matrix is converted to a quaternion using a numerically stable batched algorithm. For appearance, we compute ellipsoid-normalized directions for each point and fit spherical harmonics coefficients to the RGB values via regularized batched least-squares. Opacity is estimated by averaging and logit-transforming the per-point values.
Algorithm 1 SF-VAE: one training step for a minibatch of submanifold fields
1:Batch $\{(\bm{\mu}_{i},\bm{\Sigma}_{i},\mathbf{c}_{i},o_{i})\}_{i=1}^{B}$, fixed $r^{2}{=}1$, point count $P$
2:for each $i$ do
3:  (Sampling on $\mathcal{M}_{i}$) Sample $\{\mathbf{u}_{k}\}_{k=1}^{P}$ quasi-uniformly on $\mathbb{S}^{2}$; set $\mathbf{x}_{i,k}\leftarrow\bm{\mu}_{i}+\bm{\Sigma}_{i}^{1/2}\mathbf{u}_{k}$ so that $(\mathbf{x}_{i,k}-\bm{\mu}_{i})^{\top}\bm{\Sigma}_{i}^{-1}(\mathbf{x}_{i,k}-\bm{\mu}_{i})=r^{2}$
4:  (Color/opacity) $\mathbf{d}_{i,k}\leftarrow(\mathbf{x}_{i,k}-\bm{\mu}_{i})/\|\mathbf{x}_{i,k}-\bm{\mu}_{i}\|$;   $\mathbf{c}_{i,k}\leftarrow\mathrm{SH}(\mathbf{c}_{i},\mathbf{d}_{i,k})$;   $\alpha_{i,k}\leftarrow\sigma(o_{i})$
5:  (Point set) $\mathcal{P}_{i}\leftarrow\{(\mathbf{x}_{i,k},\mathbf{c}_{i,k},\alpha_{i,k})\}_{k=1}^{P}$
6:end for
7:(Encode) $(\bm{\mu}^{z}_{i},\log\bm{\sigma}^{2,z}_{i})\leftarrow E(\mathcal{P}_{i})$;  $\mathbf{z}_{i}\leftarrow\bm{\mu}^{z}_{i}+\bm{\sigma}^{z}_{i}\odot\varepsilon,\;\varepsilon\sim\mathcal{N}(0,I)$
8:(Decode) $(\widehat{\mathcal{M}}_{i},\widehat{F}_{i},\hat{\alpha}_{i})\leftarrow D(\mathbf{z}_{i})$;  $\widehat{\mathcal{P}}_{i}\leftarrow\{(\hat{\mathbf{x}}_{i,k},\widehat{F}_{i}(\hat{\mathbf{x}}_{i,k}),\hat{\alpha}_{i})\}$
9:(Recover parameters) $\widehat{\bm{\Sigma}}_{i}\leftarrow\mathrm{PCA}(\{\hat{\mathbf{x}}_{i,k}\})$;  $\widehat{\mathbf{c}}_{i}\leftarrow\arg\min_{\mathbf{c}}\sum_{k}\|\mathrm{SH}(\mathbf{c},\hat{\mathbf{d}}_{i,k})-\widehat{F}_{i}(\hat{\mathbf{x}}_{i,k})\|_{2}^{2}$
10:(Loss) $\mathcal{L}_{\text{rec}}\leftarrow W_{2}^{(\varepsilon)}(\mathcal{P}_{i},\widehat{\mathcal{P}}_{i})$ with $d(\cdot,\cdot)$ from Eq. (8);  $\mathcal{L}\leftarrow\mathcal{L}_{\text{rec}}+\beta\cdot\mathrm{KL}$
11:(Update) $\bm{\theta}_{E},\bm{\theta}_{D}\leftarrow\bm{\theta}_{E},\bm{\theta}_{D}-\eta\nabla_{\bm{\theta}_{E},\bm{\theta}_{D}}\mathcal{L}$
C.2 Generated 3D Gaussian primitives dataset
Parameter priors and sampling Each primitive $\mathcal{G}_{i}$ is sampled as $\bm{\theta}_{i}=\{\bm{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\mathbf{c}_{i},o_{i}\}$ and converted to $(\bm{\mu}_{i},\Sigma_{i},\mathbf{c}_{i},\alpha_{i})$ with $\Sigma_{i}=R(\mathbf{q}_{i})\,\mathrm{diag}(\exp(\mathbf{s}_{i}))^{2}\,R(\mathbf{q}_{i})^{\top}$ and $\alpha_{i}=\sigma(o_{i})$. Unless otherwise stated we use:
  • Mean. $\bm{\mu}_{i}=(0,0,0)$. Since our setting only samples single Gaussians, extrinsic information is not needed.
  • Rotation. $\mathbf{q}_{i}$ is sampled uniformly on $SO(3)$ (normalized and, if needed, enforce a canonical sign).
  • Scale. Log–axes $\mathbf{s}_{i}\in\mathbb{R}^{3}$ drawn i.i.d. from $\mathcal{U}([s_{min},s_{max}])$; set activated scales $\exp(s_{i})$.
  • SH coefficients. Let $\beta>1$ be the decay factor (in code, $\beta=4$). For degree $\ell=0,\dots,L$, we draw the $(2\ell{+}1)$-dimensional SH band as $\mathbf{c}_{\ell}\sim\mathcal{N}\!\big(\mathbf{0},\,\sigma_{\ell}^{2}I_{2\ell+1}\big),\qquad\sigma_{\ell}\;=\;\beta^{-\ell},$ i.e., $\mathrm{Var}[\,c_{\ell,m}\,]\;=\;\beta^{-2\ell}\quad\text{for }m=-\ell,\dots,\ell.$ If coefficients above the chosen degree $L$ are padded up to $L_{\max}$, we use i.i.d. noise $c_{\ell,m}\sim\mathcal{N}(0,\sigma_{\text{void}}^{2})$ with $\sigma_{\text{void}}=0.05$.
  • Opacity. Logit $o_{i}\sim\mathcal{U}([o_{min},o_{max}])$, with $\alpha_{i}=\sigma(o_{i})$.
To sum up, we can describe the data distribution in this dataset as:
$\mathcal{D}$ $=\bigl\{(\bm{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i},\mathbf{c}_{i},o_{i})\bigr\}_{i=1}^{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\underbrace{\delta_{\mathbf{0}}}_{\bm{\mu}}\times\underbrace{\mathcal{U}\!\big(SO(3)\big)}_{\mathbf{q}}\times\underbrace{\mathcal{U}\!\big([s_{\min},s_{\max}]\big)^{3}}_{\mathbf{s}}$
$\quad\times\;\underbrace{\left(\prod_{c\in\{R,G,B\}}\Big[\prod_{\ell=0}^{L}\mathcal{N}\!\big(0,\beta^{-2\ell}\big)^{2\ell+1}\;\times\;\prod_{\ell=L+1}^{L_{\max}}\mathcal{N}\!\big(0,\sigma_{\text{void}}^{2}\big)^{2\ell+1}\Big]\right)}_{\mathbf{c}\in\mathbb{R}^{3\times K},\;K=(L_{\max}+1)^{2}}$
[…]
where $\beta>1$ is the SH variance–decay factor (default $\beta{=}4$) and $\sigma_{\text{void}}$ is the padding noise std (default $0.05$). Activations used downstream are $\Sigma_{i}=R(\mathbf{q}_{i})\,\mathrm{diag}\!\big(\exp(\mathbf{s}_{i})\big)^{2}R(\mathbf{q}_{i})^{\top}$ and $\alpha_{i}=\sigma(o_{i})$.
Defaults and ablations Default hyperparameters: N=500N{=}500K, $L_{\max}{=}3$ (K=16K{=}16), P=144P{=}144, $s_{min}{=}-8$, $s_{max}{=}0$, $o_{min}{=}-5$, $o_{max}{=}10$. These parameters are built from statistical analysis of data distribution in diverse 3DGS datasets. We ablate $P\in\{36,64,144,576\}$ in the main paper to assess reconstruction vs. cost.
Data Formatting and Processing. For each of the randomly synthesized primitive we store the native tuple $\bm{\theta}_{i}$ (float32 arrays for $\bm{\mu}$, $\mathbf{q}$, $\mathbf{s}$, $\mathbf{c}$, $o$). The discretized field $\mathcal{P}_{i}$ is sampled from the primitive of $\bm{\theta}_{i}$ to a tensor of $(B,P,7)$ where $7$ is $(x,y,z,r,g,b,\alpha)$. It can be obtained at runtime to reduce memory and storage requirements. The dataset exposes toggles to return either $\bm{\theta}$ or $\mathcal{P}$.
Algorithm 2 GaussianGen: Generate one colored point cloud from a random Gaussian
1:point count $P$ (default 144144), SH degree $L_{\max}$ with $K=(L_{\max}{+}1)^{2}$
2:Sample raw parameters $\bm{\mu}\in\mathbb{R}^{3}$, $\mathbf{s}\in\mathbb{R}^{3}$, $\mathbf{q}\in SO(3)$, $o\in\mathbb{R}$, $\mathrm{feat\_dc}\in\mathbb{R}^{3}$, $\mathrm{feat\_extra}\in\mathbb{R}^{3(K-1)}$
3:Assemble SH coefficients $\mathbf{C}\in\mathbb{R}^{3\times K}$ by stacking per channel:
4:    $\mathbf{C}_{r}\leftarrow[\mathrm{feat\_dc}[0],\;\mathrm{feat\_extra}[0:(K-1)]]$
5:    $\mathbf{C}_{g}\leftarrow[\mathrm{feat\_dc}[1],\;\mathrm{feat\_extra}[(K-1):2(K-1)]]$
6:    $\mathbf{C}_{b}\leftarrow[\mathrm{feat\_dc}[2],\;\mathrm{feat\_extra}[2(K-1):3(K-1)]]$
7:Activate parameters
8:    $\mathbf{q}\leftarrow\mathbf{q}/\|\mathbf{q}\|$;  $R\leftarrow R(\mathbf{q})$;  $\bm{\sigma}\leftarrow\exp(\mathbf{s})$;  $\alpha\leftarrow\sigma(o)$
9:Build surface grid
10:    $n\leftarrow\sqrt{P}$;  create angular grids $u\!\in\![0,2\pi)$ and $v\!\in\![0,\pi]$ of size $n\times n$
11:    directions $\mathbf{d}(u,v)\in\mathbb{S}^{2}$ from spherical angles $(u,v)$
12:Map to ellipsoid (iso-density $r^{2}{=}1$)
13:    $\mathbf{x}(u,v)\leftarrow\bm{\mu}+R\,\mathrm{diag}(\bm{\sigma})\,\mathbf{d}(u,v)$
14:Evaluate color field
15:    $\mathbf{c}(u,v)\leftarrow\mathrm{SH}(\mathbf{C},\,\mathbf{d}(u,v))$
16:    (optional) color post-process: $\mathbf{c}\leftarrow\mathrm{clip}(\mathbf{c}{+}0.5,\;0,1)$
17:Pack output
18:    replicate $\alpha$ to all points;  flatten grids to $P$ points
19:    return $\{(\mathbf{x}_{k},\;\mathbf{c}_{k},\;\alpha)\}_{k=1}^{P}\in\mathbb{R}^{P\times 7}$   // $(x,y,z,r,g,b,\alpha)$
Appendix D Additional Evaluation and Visualization Results
Figure 7: Sample visual comparison of linear interpolation in submanifold field embedding space and parametric space. Interpolation in SF embedding shows smooth transition from source to target. Perturb $\mathbf{s}$ and $\mathbf{q}$ will dramatically change the geometry of the resulted Gaussian, verifying the feature heterogeneous problem.
Figure 8: When rotation $\mathbf{q}$ is inverted to $\mathbf{q}$, the submanifold field model (left) can still correctly reconstruct the scene, while parametric model fails to process this equivalent rotation.
Figure 9: Full qualitative results for rasterization using different model configurations on Mip-NeRF 360.

### v2

@@
 In this paper, we propose an embedding representation of 3DGS based on continuous submanifold fields that encapsulate the intrinsic information of Gaussian primitives, thereby benefiting the learning of 3DGS.
+Implementation available at https://github.com/cilix-ai/gs-embedding.
 1 Introduction
@@
 Despite its effectiveness in optimization-based reconstruction, we identify fundamental limitations when this parametric representation is employed as a learning space for neural networks.
-Our analysis reveals three critical issues.
-First, the mapping from parameters to rendered output exhibits non-uniqueness: multiple distinct parameter configurations yield identical visual results.
-This occurs through quaternion sign ambiguity (i.e., $\mathbf{q}$ and $\mathbf{q}$ represent identical rotations), symmetry-induced rotation invariance, and equivalences in spherical harmonic coefficients.
-Second, the parameter components exhibit severe numerical heterogeneity, with scales spanning large magnitude while quaternions remain unit-normalized.
-Third, these parameters inherently reside on distinct mathematical manifolds: positions in $\mathbb{R}^{3}$, rotations in $\text{SO}(3)$, and appearance in spherical harmonic space, inconsistent with the Euclidean space demands of standard neural architectures.
-These theoretical limitations translate into substantial practical failures.
-Our empirical analysis reveals critical instabilities: negating quaternion values, which should produce equivalent rotations, causes complete decoding failure in parameter-trained autoencoders (see App.
+Specifically, the native parameterization $\bm{\theta}$ conflicts with the inductive biases of standard neural architectures in three critical ways.
+First, the mapping from parameters to rendered output exhibits non-uniqueness.
+Ambiguities such as quaternion sign duality and symmetry-induced variances create a one-to-many mapping.
+This mapping creates “embedding collisions” where distinct parameter inputs producing identical visual outputs generate conflicting supervision signals, leading to training instability and poor convergence (Bengio et al., 2013; Wang and Isola, 2020).
+Second, the parameter components suffer from numerical heterogeneity.
+Spatial positions span large magnitudes while quaternions remain unit-normalized, violating the homogeneous feature distribution assumption required for effective gradient flow (Ioffe and Szegedy, 2015).
+Third, these parameters inherently reside on distinct mathematical manifolds, such as positions in $\mathbb{R}^{3}$, rotations in $SO(3)$, and appearance in spherical harmonic coefficients.
+Forcing these non-Euclidean variables into the Euclidean feature spaces of standard encoders breaks their intrinsic geometric structure, making the representation difficult to compress or regularize effectively.
+These theoretical misalignments translate into substantial practical failures across various applications.
+Our empirical analysis reveals that parametric encoders are critically unstable; for instance, due to the sign ambiguity, simply negating a quaternion (a mathematically equivalent rotation) can cause complete decoding failure in parameter-trained autoencoders, see App.
 Fig.
-8).
-Parametric representations achieve lower reconstruction quality and exhibit severe degradation under domain shift, particularly when transferring from object-level to scene-level data.
-Moreover, parametric embeddings lack robustness to noise, with minor perturbations causing disproportionate reconstruction errors.
-These consistent failures indicate fundamental representation deficiencies.
-We thus propose a principled alternative: representing each Gaussian primitive as a continuous field defined on its iso-probability surface.
-This submanifold field representation establishes a unique relationship between Gaussians and their geometric-photometric properties, eliminating the ambiguity inherent in parametric representations.
-By discretizing this field as a colored point cloud sampled from the probability surface, we obtain a representation that maintains uniform numerical properties and respects the underlying geometric structure.
-Particularly, we employ a variational autoencoder to learn embeddings from discretized submanifold fields, introducing a Manifold Distance metric based on optimal transport that better aligns with perceptual quality than parameter-space distances.
-Extensive experiments demonstrate substantial improvements in reconstruction quality, superior cross-domain generalization, and more robust latent representations compared to parametric baselines.
-The learned embeddings exhibit better noise resilience and semantic structure, validating our submanifold field representation as a more suitable learning space for 3D Gaussian primitives.
+9.
+This instability extends to generative tasks involving latent manipulation, where we observe that linear interpolation in the parametric latent space results in geometric “jitters” and unnatural transitions.
+Furthermore, parametric embeddings lack robustness to noise, with minor perturbations causing disproportionate reconstruction errors.
+Crucially, for downstream tasks extensively explored in recent work including generative modeling (Yi et al., 2024; Zhou et al., 2024b), compression (Shin et al., 2025; Girish et al., 2024), and editing (Chen et al., 2024a), these flaws manifest as discontinuous latent spaces and inefficient encoding.
+Instead of capturing robust geometric semantics, models are forced to resolve parameter ambiguities with more efforts, leading to sub-optimal performance in learning-based reconstruction (Charatan et al., 2024) and limited generalization across diverse domains.
+We thus propose a principled alternative that represents each Gaussian primitive as a continuous field defined on its iso-probability surface.
+This submanifold field representation establishes a unique correspondence between Gaussians and their geometric-photometric properties, removing the ambiguities of parametric representations.
+By discretizing this field as a colored point cloud sampled from the probability surface, we obtain a numerically uniform and geometrically consistent representation.
+We further employ a variational autoencoder to learn embeddings from these discretized submanifold fields, together with a Manifold Distance metric based on optimal transport that better correlates with perceptual quality than parameter-space distances.
+Extensive experiments show higher reconstruction quality, stronger cross-domain generalization, and more robust latent representations.
+The learned embedding space is smooth and generalizable, indicating strong potential for semantic understanding and generative modeling, as demonstrated through unsupervised segmentation and neural field decoding with the proposed embeddings.
 To summarize the contribution of this work, we:
@@
 (2024) generate pixel-wise parameter maps and lift them to 3D via depth estimation.
-This paradigm has been extended to pose-free settings (Hong et al., 2024; Chen et al., 2024c; Tian et al., 2025), and transformer-based methods further improve scalability and generalization (Li et al., 2025b; Jiang et al., 2025; Lin et al., 2025b).Overall, these methods directly output Gaussian parameters from neural networks, showing that 3DGS can serve as an effective target for network-based prediction in efficient reconstruction.
+This paradigm has been extended to pose-free settings (Hong et al., 2024; Chen et al., 2024c; Tian et al., 2025), and transformer-based methods further improve scalability and generalization (Li et al., 2025b; Jiang et al., 2025; Lin et al., 2025b).
+Overall, these methods directly output Gaussian parameters from neural networks, showing that 3DGS can serve as an effective target for network-based prediction in efficient reconstruction.
 Embedding Gaussian Primitives.
@@
 (2024) learn structured latent variables from 3D Gaussian space to fulfill generation tasks.
-Editing and style-transfer methods use diffusion or style conditioning to manipulate Gaussians primitives in latent or rendering spaces (Chen et al., 2024a; Vachha & Haque, 2024; Lee et al., 2025; Palandra et al., 2024; Zhang et al., 2024; Kovács et al., 2024; Yu et al., 2024).
+Editing and style-transfer methods use diffusion or style conditioning to manipulate Gaussians primitives in latent or rendering spaces (Chen et al., 2024a; Vachha and Haque, 2024; Lee et al., 2025; Palandra et al., 2024; Zhang et al., 2024; Kovács et al., 2024; Yu et al., 2024).
 Other works improve rendering quality by optimizing Gaussian parameters under diffusion priors (Tang et al., 2023; Yi et al., 2024; Chen et al., 2024d), while compression methods (Girish et al., 2024; Yang et al., 2025) reduce storage and computation by quantizing and embedding Gaussian parameters.
@@
 Formally, there exist at least two distinct parameter sets, $\bm{\theta}_{1}\in\Theta$ and $\bm{\theta}_{2}\in\Theta$ with $\bm{\theta}_{1}\neq\bm{\theta}_{2}$, that generate the exact same field $\phi_{\mathcal{G}}\in\Phi$.
-The non-uniqueness is from quaternion sign ambiguity, geometric symmetries, and rotation-spherical harmonic interactions producing equivalent parameter combinations (see Appendix for proof).
-The non-uniqueness of $\bm{\theta}$ will create “embedding collisions” where different parameter vectors produce identical rendered output (Wang & Isola, 2020).
+The non-uniqueness is from quaternion sign ambiguity, geometric symmetries, and rotation-spherical harmonic interactions producing equivalent parameter combinations (see App.
+A for proof).
+The non-uniqueness of $\bm{\theta}$ will create “embedding collisions” where different parameter vectors produce identical rendered output (Wang and Isola, 2020).
 This makes the learning objective $\|\bm{\theta}_{\text{pred}}-\bm{\theta}_{\text{target}}\|_{p}$ ambiguous, as multiple parameter configurations can achieve the same visual result.
@@
 Numerical Heterogeneity The parameter components violate the homogeneous distribution assumption of standard neural architectures.
-Neural networks typically assume features share similar statistical properties for effective gradient flow (Ioffe & Szegedy, 2015).
+Neural networks typically assume features share similar statistical properties for effective gradient flow (Ioffe and Szegedy, 2015).
 However, 3D Gaussian parameters span vastly different ranges.
@@
 This setting further decouples evaluation results with the training objective functions.
-Apart from input dimension, all models share identical depth, width, latent size (32), and optimizer settings, ensuring a matched capacity.
+Apart from input dimension, all models share identical depth, width, latent size (32), and optimizer settings (all using Adam), ensuring a matched capacity.
 Datasets.
@@
 3
-4.2 Quantitative Reconstruction Results
+4.2 Evaluation on Representation Learning Framework
 Table 1: Reconstruction quality comparison for object-level (ShapeSplat) and scene-level (Mip-NeRF 360) datasets.
@@
 Input Representation Encoder Decoder PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ M-Dist $L_{1}$-Dist
-[…]
+ShapeSplat
 Parametric MLP MLP 37.512 0.888 0.152 0.184 0.040
@@
 Submanifold Field SF-VAE SF-VAE 63.408 0.990 0.010 0.041 0.098
-[…]
+Mip-NeRF 360
 Parametric MLP MLP 18.818 0.564 0.452 0.510 0.034
@@
 Submanifold Field SF-VAE SF-VAE 29.833 0.953 0.079 0.048 0.179
+Zero-shot Reconstruction.
 We present a comprehensive quantitative and qualitative analysis of reconstruction quality for both object-level and scene-level data, as summarized in Tab.
-4.2 and Fig.
+1 and Fig.
 3.
@@
 Similar performance gains are observed in scene-level reconstruction, where the submanifold field model demonstrates better performance across diverse spatial contexts.
-Importantly, the Manifold Distance (M-Dist) metric shows a stronger empirical correlation with perceptual metrics like PSNR and LPIPS than traditional $L_{1}$ parameter distances, supporting our claim that M-Dist is a more robust and meaningful similarity measure for 3D Gaussian representations, truthfully reflecting perceptual differences rather than merely parameter discrepancies.
+Importantly, the Manifold Distance (M-Dist) metric shows a stronger empirical correlation with quality metrics like PSNR and LPIPS than traditional $L_{1}$ parameter distances, supporting our claim that M-Dist is a more robust and meaningful similarity measure for 3D Gaussian representations, truthfully reflecting perceptual differences rather than merely parameter discrepancies.
 The consistent improvement margin across both datasets highlights the advantage of learning in the submanifold field space, which better preserves intrinsic structure and view-dependent appearance, confirming the efficacy of our representation for high-fidelity 3D Gaussian modeling.
@@
 Samples selected arbitrarily from Mip-NeRF 360 and ShapeSplat.
-Parametric models can induce confusion in rotation space, failing to predict the correct rotation.
-4.3 Cross-Domain Generalization
-Table 2: Results on representation generalization: train on one dataset, test on the other.
+Parametric models can induce confusion in parameter space, failing to embed and restore the correct Gaussian parameters.
+Table 2: Reconstruction quality comparison under cross-domain setting.
+All models trained on either ShapeSplat or Mip-NeRF 360 dataset are tested on another dataset.
+We show that the generalization ability of SF Embedding framework is inherently domain-agnostic even without random data.
 Train set Test set Input Represent. Encoder / Decoder PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
 Submanifold Field SF-VAE / SF-VAE 62.576 0.990 0.014
-We evaluate cross-domain generalization by training on one dataset and testing on the other (object-level $\leftrightarrow$ scene-level) under an identical training protocol and capacity budget as in the reconstruction study.
-Concretely, we train either the proposed SF-VAE or the parametric MLP baselines on a source set and evaluate zero-shot on a target set, rendering novel views and reporting PSNR, SSIM, and LPIPS averaged over test samples (see Tab.
+Cross-domain Reconstruction.
+We also evaluate cross-domain generalization by training on one real-world dataset and testing on the other (object-level $\leftrightarrow$ scene-level) under an identical training protocol and capacity budget as in the reconstruction study.
+Concretely, we train either the proposed SF-VAE or the parametric MLP baselines on a source set and evaluate on a target set, rendering novel views and reporting PSNR, SSIM, and LPIPS averaged over test samples (see Tab.
 2).
-Across both transfer directions, the submanifold field model consistently achieves higher reconstruction quality than the parametric baseline, with the largest margins when transferring from object-level sources to larger-scale scenes, indicating reduced sensitivity to dataset-specific statistics (scale, lighting, SH complexity).
-4.4 Latent Space Evaluation
+Across both transfer directions, the SF-based embedding consistently achieves higher reconstruction quality than the parametric baseline, indicating reduced sensitivity to domain-specific statistics (e.g., scale, lighting, SH complexity).
+Particularly, comparing these transfer results with Tab.
+1, we find that the model trained on synthetic random data actually outperforms the models transferred from real-world domains.
+This indicates that our random learning strategy effectively strips away domain priors, which establishes our approach as a unified representation, where a single model trained on synthetic data can fundamentally generalize to real-world contexts.
+4.3 Sensitivity Study of Representation
 Robustness to Noise.
-To evaluate the submanifold field embedding’s robustness to noise, we gradually add higher levels of gaussian noise to the embedding space of each representation and test their reconstruction quality and M-Dist.
+To evaluate the submanifold field embedding’s robustness to noise, we gradually add higher levels of gaussian noise to the embedding space of the parametric model and the submanifold field model and test their reconstruction quality and M-Dist.
+To ensure fair comparison, we use noise level as a ratio to variance instead of absolute noise magnitude.
 As shown in Fig.
-4, the embedding space of submanifold field model is more robust to random perturbation, this makes submanifold field embeddings a better learning target since it is less sensitive to potential systematic noise.
-Figure 4: Parametric results use MLP encoder and SF-VAE’s decoder.
-Left: Sample visual comparison on robustness to noise in embedding space.
+4, the embedding space of submanifold field model is more robust to random perturbation, this makes submanifold field embeddings a better learning target since it is less sensitive to potential noise introduced by downstream regression.
+Figure 4: Reconstruction results using embeddings with noise.
+Left: Visualization of reconstructed scene from noisy embeddings of Gaussian parameters (MLP) and SF-VAE.
 Right: Comparison on M-Dist for different noise levels added to embedding space, tested on Mip-NeRF 360.
-Latent Space Clustering.
+Noise level is defined as the ratio between the noise magnitude and the embedding variance.
+Latent Space Interpolation.
+To evaluate the regularity of the latent space of the proposed representation, we randomly sample pairs of source and target Gaussian primitives $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ and linearly interpolate each pair for a fixed number of steps $n=7$.
+Compared with parametric space, the interpolation in submanifold field embedding space shows a smooth transition path, while interpolation in parametric space shows undesired jitter in rotation and scale, indicating space irregularities, see App.
+D.
+This highlights the motivation to learn in the unified submanifold field embedding space.
+4.4 Representation Applicability
+Unsupervised Clustering.
 To further probe the semantic structure of the learned embedding spaces, we perform unsupervised graph clustering on both the raw Gaussian parameter space and the embedding outputs of each model.
@@
 This indicates that the submanifold field embedding captures more dense semantics and discriminative features, validating its usefulness.
-Figure 5: Qualitative evaluation of unsupervised graph clustering, compared using Gaussian parameters and the embedding of the respective models, using the same hyperparameters.
-Submanifold field embeddings show better preservation of detailed semantics, validating its usefulness.
-Latent Space Interpolation.
-To evaluate the regularity of the latent space of the proposed representation, we randomly sample pairs of source and target Gaussian primitives $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ and linearly interpolate each pair for a fixed number of steps $n=7$.
-Compared with parametric space, the interpolation in submanifold field embedding space shows a smooth transition path, while interpolation in parametric space shows undesired jitter in rotation and scale, indicating space irregularities, see App.
-D.
-This highlights the motivation to learn in the unified submanifold field embedding space.
-4.5 More Studies
-Figure 6: Ablation study results, tested on Mip-NeRF 360.
-From left to right: (a) embedding space dimension, (b) generated training dataset size, (c) Submanifold Field sample size (all values normalized to $[0,1]$, with LPIPS inversely normalized).
-Latent Space Dimension.
+Figure 5: Unsupervised graph clustering based on raw Gaussian parameters and various embeddings.
+Submanifold field embeddings show better preservation of detailed semantics, showing its downstream applicability.
+Gaussian Neural Fields.
+To validate the potential of our representation for advanced downstream tasks, we introduce the Gaussian Neural Field (GNF).
+Drawing inspiration from the decoding structures in generative diffusion models (e.g., DiffGS by Zhou et al.
+(2024b)) and neural compression frameworks (Wu and Tuytelaars, 2024), the GNF functions as a coordinate-based neural implicit field as illustrated in Fig.
+6.
+Specifically, it employs a lightweight MLP (architecture detailed in App.
+D.4) to learn a continuous mapping from spatial coordinates $\mathbf{x}_{i}$ to per-primitive descriptors.
+This setup allows us to evaluate the “learnability” of our representation: while regressing heterogeneous raw parameters $\bm{\theta}_{i}$ often leads to optimization difficulties, our unified SF embeddings provide a smooth and well-conditioned target for the neural field.
+As evidenced in Tab.
+3 and visualization in App.
+D.4, the SF-guided GNF outperforms the parameter-based baseline in visual fidelity with equivalent training effort.
+This indicates that our representation is more friendly to neural networks, hinting at its utility for potential downstream generative and compression tasks.
+Figure 6: Setting of a Gaussian Neural Field, we compare between the prediction target SF embedding and raw GS parameters.
+Table 3: Comparison between Gaussian Neural Fields trained using submanifold field embeddings and raw Gaussian parameters. Top: ShapeSplat, bottom: Mip-NeRF 360.
+Target PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # Params
+Raw GS Parameter 51.660 0.925 0.141 0.21M
+SF Embedding 58.619 0.980 0.043 0.20M
+Raw GS Parameter 19.922 0.648 0.410 1.87M
+SF Embedding 24.395 0.804 0.261 1.85M
+4.5 More Studies on Implementation Details
+Figure 7: Behavior studies tested on Mip-NeRF 360. From left to right: (a) embedding dimension, (b) generated training dataset size, (c) Submanifold Field discretized (i.e., point sample) grid size.
+[…]
+Behavior Study of Latent Space Dimension.
 We evaluated different embedding space dimensions for the SF-VAE model to meet the best trade-off between compression and reconstruction quality.
@@
 4.3 suggests potential in further latent tokenization and quantization.
-Size of Training Set.
-To empirically determine the number of random training samples required to achieve the best reconstruction results, we tested a wide spectrum of training sample sizes, ranging from 5K to 500K (baseline).
+Behavior Study of Training Set Size.
+To determine the number of random training samples required to achieve the best reconstruction results, we vary the trainiing sample size from 5K to 500K (baseline), see Fig.
+7 (b).
 The results indicate the proposed representation is data efficient.
 When only using 2% of the baseline training sample, our model can achieve close-to-baseline performance.
-Point Sample Size of Submanifold Fields.
-To ensure the submanifold fields are truthfully represented in a discrete manner, we evaluate different sample sizes $P$.
+Behavior Study of SF Discretized Size.
+To ensure the submanifold fields are truthfully represented in a discrete manner, we evaluate different sample sizes $P$, see Fig.
+7 (c).
 Going from the lowest tested $P=6^{2}$ to the baseline $P=12^{2}$, we observe a gradual improvement in reconstruction quality, while going above $P=12^{2}$ yields very little improvement.
@@
 Computational Efficiency.
-Increasing the point sample size $P$ leads to higher compute time and memory consumption.
-Encoding time stays low since a lightweight PointNet-based encoder shares weights with all input points, ensuring efficient encoding of large-scale scenes, with an inference speed of 1.72s per 1 million Gaussians for $P=12^{2}$ using a batch size of 4096 on an RTX 5090.
-Query time during decoding is $O(P)$ which is $O(n^{2})$ w.r.t.
+Increasing the point sample size $P$ increases computation and memory.
+Encoding time remains low since a lightweight PointNet-based encoder shares weights with all input points, giving an inference speed of 1.72s per 1 million Gaussians for $P=12^{2}$ with a batch size of 4096 on an RTX 5090.
+Decoding time is 4.20s per 1 million Gaussians, from embedding to Gaussian parameters.
+The complexity is $O(P)$ or $O(n^{2})$ w.r.t.
 the grid size $n$.
-When the computation is not bottlenecked and uses a large enough batch size, the amortized time complexity is $\Theta(n)$.
+We utilize the advantage of GPU parallel computation to boost the calculation for parameter fitting module.
+In the 4.20s of decoding time, the fitting module (PCA + SH fitting) consumes only about 0.48s which is negligible for large Gaussian scenes.
 5 Conclusion and Limitations
@@
 Our current setup operates at the single-Gaussian level, while this ensures data invariance, it omits explicit inter-splat structure modeling for more complex representation learning.
-Promising directions include set/scene-level encoders with permutation-invariant attention, point cloud to 3DGS inpainting, generative modeling with submanifold field embeddings, temporal extensions to dynamic scenes, and applications to compression, retrieval, and regularization in broader 3DGS pipelines.
+Promising directions include set/scene-level encoders with permutation-invariant attention, point cloud to 3DGS inpainting, generative modeling with submanifold field embeddings, temporal extensions to dynamic scenes, and applications to compression, retrieval, and regularization in broader downstream 3DGS pipelines.
 Appendix A Proof of Proposition 1
@@
 Appendix D Additional Evaluation and Visualization Results
+D.1 Interpolation & Noise Visualizations
 Figure 8: Sample visual comparison of linear interpolation in submanifold field embedding space and parametric space.
@@
 Figure 9: When rotation $\mathbf{q}$ is inverted to $\mathbf{q}$, the submanifold field model (left) can still correctly reconstruct the scene, while parametric model fails to process this equivalent rotation.
+D.2 Additional Reconstruction Visualizations
 Figure 10: Full qualitative results for rasterization using different model configurations on Mip-NeRF 360.
+D.3 Unsupervised Clustering Quantitative Results
+Clustered Feature Silhouette $\uparrow$ Cluster Compactness (MSE $\downarrow$)
+Parameters 0.090 71.306
+MLP / MLP 0.084 50.761
+MLP / SF-VAE 0.086 50.916
+SF-VAE 0.113 48.282
+Table 4: Quantitative comparisons on unsupervised clustering based on the respective clustered feature. Higher silhouette is better; lower compactness (MSE) is better.
+D.4 Gaussian Neural Field Details & Visualizations
+Network Structure for Gaussian Neural Fields.
+We define a compact per-scene MLP that takes only 3D coordinates ((x,y,z)), applies sinusoidal/Fourier positional encoding, and passes the encoded vector through a fully connected backbone (256-width 4 layers for ShapeSplat, 512-width 8 layers for Mip-NeRF 360) with a single mid-network skip that re-concatenates the encoded input for better conditioning.
+The same network architecture is used for both targets; the only difference is the final output dimension: either 32 for the SF embedding or 56 for the no-position Gaussian parameters ($q(4),s(3),\text{SH}(3K),o(1)]$).
+Hidden layers use standard ReLU/SiLU activations, the output is linear, and the overall footprint remains lightweight ($\approx 2\times 10^{5}$ parameters for ShapeSplat, $\approx 1.8\times 10^{6}$ for Mip-NeRF 360) for querying at Gaussian centers to yield SF embeddings or directly renderable Gaussian parameters.
+Figure 11: Arbitrarily picked visual comparisons of Gaussian Neural Fields trained with target Gaussian parameters and submanifold field embeddings.
+The submanifold field embedding results is qualitatively better under matching settings, indicating that the SF embedding is a better learning target for regression tasks such as training a neural field.